# SoK: A Literature and Engineering Review of Regular Expression Denial of Service (ReDoS)

### v3

@@ 4.1.6. Other ReDoS-Safe Engines
Two major programming languages, Rust and Go, have regex engines that were developed with ReDoS safety in mind (Rus, 2024a; Gor, [n. d.]). These engines use an explicit automaton representation that is simulated with Thompson’s algorithm (Cox, 2007). They also do not support E-regex features. They, therefore, offer linear-time match guarantees in the length of the input and the automaton. Turoňová observed that the construction of the automaton might be slow or the memory footprint problematic if the regex uses counters (i.e., bounded quantifiers) (Turoňová et al., 2022).
Although not included in our analysis, independent third-party regex engines such as RE2 (Google) (Goo, 2024) and Hyperscan (Intel) (Wang et al., 2019; Wang, [n. d.]) are well-known for their ReDoS resilience. RE2, a Thompson-style engine, incorporates DFA state caching optimizations (Moseley et al., 2023). While generally ReDoS-safe and non-backtracking, RE2 can, in rare cases, exhibit backtracking behavior (Davis et al., 2021). Similarly, Hyperscan is another efficient, ReDoS-safe engine, utilizing hardware acceleration alongside Glushkov’s NFA construction (Glushkov, 1961). Neither engine supports E-regex features.
In conclusion, these developments in regex engines demonstrate significant progress in mitigating ReDoS risks but also highlight ongoing challenges in balancing the trade-off between safety and usability. Modern regex engines mitigate ReDoS with non-backtracking algorithms like memoization, Thompson-style simulation, or Brzozowski derivatives to assure linear-time matching. However, the benefit comes at the cost of reduced expressiveness. Non-backtracking engines lack support E-regex features. Defenses often supplement with timeouts or fallback mechanisms for E-regexes to maintain compatibility with existing codebases.
@@ 4.2. Updated Measure of ReDoS Vulnerabilities
We answered this question through measurements on a sample from the Davis et al.’s (Davis et al., 2019) polyglot regex corpus, comprising $\sim$6K SL-regexes extracted from open-source repositories across many programming languages. We filtered this corpus for all regexes predicted to be SL by vuln-regex-detector (dav, 2024), a tool developed by Davis et al. (Davis et al., 2018a), which relies on (Rathnayake and Thielecke, 2017; Wüstholz et al., 2017; Weideman et al., 2016; Shen et al., 2018). We evaluated a sample of 500 exponential-time regex candidates (out of 3K examples) and 500 polynomial-time regex candidates (out of $\sim$100K examples). For these regexes, we generated attack inputs using, again, the vuln-regex-detector tool.
@@
Figure 5 shows our results. For each engine, we depict the number of regexes that exhibited SL behavior (as a fraction of the predicted input) in the pre- and post-ReDoS mitigation versions. Not all regexes were supported in all engines, so the candidate pool size is indicated for each engine. Observe that for the engines with mitigations, there is a substantial decrease in the number of viable SL-regexes—in other words, these mitigations appear to be effective in addressing the common causes of SL behavior. Notably, exponential behavior is resolved in C# and JavaScript but persists in Ruby and Java, while polynomial behavior continues to manifest in JavaScript and Java.
Figure 5. Performance of regex matching in old vs. latest engines in different programming languages.
@@ 5. Discussion
In this section, we distill and contextualize the main findings of our study, integrating insights from our literature review (Section 3) and engineering review (Section 4). For each idea, we discuss concrete future work directions for the community to explore.
@@
(1) ReDoS is a peculiar vulnerability. Different from more traditional vulnerabilities, e.g., SQL injections, ReDoS stands out as a contextual issue, more similar to micro-architectural attacks, where exploitability and impact are heavily dependent on the underlying system’s architecture and deployment conditions. For example, a regex pattern vulnerable to catastrophic backtracking in Node.js v15 is rendered safe in v20 due to advancements in regex engines (Section 4). When detecting ReDoS vulnerabilities, it is not enough to analyze the code of a web application, without considering its deployment configurations. For future static analysis work on ReDoS, e.g., that targets multiple languages (Hassan et al., 2023), we recommend that they consider and model the language-, engine-, and platform-specific details and differences with respect to ReDoS. For example, a super-linear regular expression might be more serious in a single-threaded web application architecture like Node.js than in inherently parallel ones like Go. For dynamic and hybrid analyses, we recommend a system perspective, similar to Shan et al. (Shan et al., 2017), in which the code under test is deployed under realistic conditions, where potential defenses like web application firewalls are also considered.
@@
(2) Strong assumptions about attackers. We find that prior work often makes strong assumptions about attackers (Section 3.2), e.g., they can inject arbitrarily large payloads, or they can control the input to any regular expression. Also, many papers neglect to deploy defenses or rely on simple rules of thumb such as that a regular expression match should not take longer than one second (Section 3.3). We think that future work should challenge these assumptions by performing empirical studies with real systems, e.g., to analyze how many regular expressions can be reached by attacker-controlled inputs. We also need more empirical studies to help the community understand how modern applications are deployed in the wild, e.g., in which specific cloud configurations and what defenses they use. Finally, future ReDoS research should perform realistic evaluations by measuring the broader impact of attacks, including system slowdowns and user experience. To aid with this, in Appendix C, we propose wrk-redos, a tool for simulating realistic ReDoS scenarios by limiting attackers’ capabilities such as the request’s length and attacks’ bandwidth. By evaluating vulnerabilities under practical conditions, tools like wrk-redos can help developers understand risks more effectively and implement appropriate mitigations.
@@
(3) Developers’ push back. To the best of our knowledge, there has been no public description of a denial-of-service attack that exploited ReDoS, so the risk posed by these vulnerabilities is still unclear, beyond academic measurement studies (Staicu and Pradel, 2018; Barlas et al., 2022). Adding to the confusion, only a small fraction of the papers we analyzed (Section 3) discuss such attacks, often in unrealistic setups. We believe that these unclear risks together with problematic security assumptions are the culprits for ReDoS vulnerabilities being a contentious topic among developers. Snyk has stated that ReDoS is the most neglected vulnerability type (ope, [n. d.]), implying that developers often ignore such findings. Also, engineers at GitHub consider it a low severity vulnerability, “not particularly serious, but easy to create by accident, obscure to understand, and sometimes tricky to fix” (Backhouse, 2023). More strident voices argue that ReDoS vulnerabilities are “malicious noise” and just lead to security fatigue (Woodruff, 2022). Several ReDoS CVE advisories have been disputed by the maintainers (e.g., CVE-2023-39663, CVE-2022-42969, CVE-2019-11391), indicating disagreement between maintainers and reporters, also reflected by negative comments in GitHub issues (Section D.4).
@@
The blame for this perspective may lie with the cybersecurity academic community. As noted in Section 3, researchers have performed large-scale measurements of ReDoS as well as produced new and more effective detectors for vulnerable regexes. Their measurement instruments have been integrated into popular security scanners, e.g., the integration of safe-regex into eslint. However, the resulting scanners analyze only the regex, without considering reachability constraints as well as other input constraints. When academics apply these tools, they (should) filter false positives based on the regex’s context. When practitioners do so, they may make much ado about nothing. This suggests that empirical studies of academic tools uptake is of interest, as well as prompting the research community to reflect on direct access to research tooling and more developer-friendly releases. More importantly, future work in the field should perform user studies with developers to understand their pain points with respect to ReDoS (tools), security assumptions, and deployment configurations. In this way, we can better align the community’s assumptions with those of practitioners.
@@
(4) Engines have evolved significantly. It is encouraging that modern language runtimes changed their implementation in response to prior work on ReDoS (Section 4). These changes were often driven or informed by researchers, e.g., Ruby’s adoption of Davis’s proposals (Davis, 2019; Davis et al., 2021). [^4] See https://bugs.ruby-lang.org/issues/19104. We believe that this is a success story like few in our community that shows how researchers and practitioners can collaborate effectively. Our study shows that modern regex engines have evolved significantly, incorporating advanced techniques such as Thompson’s algorithm, memoization, and resource capping to mitigate ReDoS. However, these solutions are often complex and rarely as clean as their theoretical counterparts. Additionally, there are often redundant implementations that are triggered on demand by the engine, when matching a regular expression. Hence, future work should investigate the deployed fixes to find potential bugs or misbehaving corner cases. We advocate for comparative studies like ours Section 4, in which multiple engines are investigated on the same regex/dataset to study the advantages and disadvantages of the deployed defenses. The research community should also propose automatic solutions for migrating to patched engines. More importantly, research on ReDoS should consider these advancements to ensure findings remain relevant to contemporary systems.
@@
(5) Performance bugs vs. security vulnerabilities. Many of our findings also apply to other algorithmic complexity vulnerabilities, e.g., Zip bombs. We note that while the possibility of weaponizing slow computation is real, attackers still seem to prefer brute-force, volume-based DDoS attacks. Researchers should aim to understand why that is the case and, simultaneously, continue to develop defenses against such potential future attacks. We also believe that there are many insufficiently-explored, cross-disciplinary opportunities for ideas that aim to bridge the gap between performance, measurement, and security communities. For example, we argue that most performance bugs, e.g., redundant loop traversals (Olivo et al., 2015b), can probably be lifted to a DoS attack when malicious intent is assumed. Thus, future work, should investigate in more detail the relation between performance bugs and CPU-based DoS attacks, in particular, in modern web applications.
@@ 6. Conclusion
In this work, we perform a multi-faceted study of the recent academic work on ReDoS and the new engineering realities in the mainstream programming languages, after initial fixes were deployed in most of them. We report on concrete examples of language runtimes that adopted defenses, which appear to be motivated by academic research. When studying the efficacy of these defenses, we find that they mitigate most ReDoS payloads, but they still leave significant room for attackers to maneuver. By surveying the academic work on ReDoS, we find that many papers in this domain use weak definitions and strong threat models, often considering any slowdown caused by regular expression matching as a security vulnerability. We advocate for future work to consider the recently-deployed fixes in mainstream programming languages and to evaluate ReDoS attacks under realistic deployment conditions, e.g., against a web application deployed in the cloud using reversed proxies and redundancy. We also propose restricting the attackers’ capabilities to realistic payload sizes and modest bandwidths that comply with the asymmetric DoS setting. More importantly, future work should study the more general problem of weaponizing slow computation, which the attackers might leverage in the near future.
7. Ethics Considerations
In our judgment, this work presents an acceptable ethical risk-reward tradeoff. Specifically:
  • The literature review (Section 3) and engineering review (Section 4) describe and systematize only publicly-available resources.
  • Any measurement tool such as wrk-redos (Section C.1) poses dual use dilemmas: it helps improve systems but may also aid attackers. We believe its benefits outweigh the risks.
  • In our evaluation of wrk-redos (Section C.2), we only rely on known/patched vulnerabilities, and we only simulate ReDoS attacks against web servers that we control.
8. Compliance with Open Science Policy
An artifact accompanies this paper at https://anonymous.4open.science/r/wrk-DoS-BD74. The artifact includes (1) the dataset of super-linear regexes with their corresponding attack strings, and (2) the source code and usage manual for the realistic ReDoS attack simulation framework proposed in Appendix C.
@@ Appendix A Regex Feature Reference
Table 6. Full set of features and notation of regexes, compiled from (Hazel, 1997). $R$ is a sub-pattern.
Abbrev. Feature Notation
@@
[…]
@@
ADD One-or-more rep. $R+$
QST Zero-or-one rep. $R?$
@@
LWB At-least-$m$ rep. $R\{m,\}$
SNG Exactly-$n$ rep. $R\{n\}$
@@
CCC Custom char. class $[aeiou]$
RNG Char. range $[a-z]$
NCCC Negated class [^$[\textasciicircum{aeiou}]$
ANY Any char. .
WSP Whitespace $\setminus{}s$
DEC Numeric $\setminus{}d$
WRD Word $\setminus{}w$
NWSP Non-whitespace $\setminus{}S$
NDEC Non-numeric $\setminus{}D$
NWRD Non-word $\setminus{}W$
VWSP Vertical space $\setminus{}v$
@@
STR Start-of-string/line ^$\textasciicircum R$, $\setminus{}\text{A}R$
@@
[…]
PLA Pos. lookahead (?$=R)$
NLA Neg. lookahead (?$!R)$
PLB Pos. lookbehind (?<${\textless}=R)$
NLB Neg. lookbehind (?<${\textless}!R)$
@@
BKR Numeric backref. $(R)\ldots\setminus{}1$
[…]
@@
LZY Non-greedy rep. $R*?,R+?,R\{m,n\}?$
ATM Atomic group (?>${\textgreater}R)$
POS Possessive quantifier $R++,R*+$

### v4

@@ 4.1.6. Other ReDoS-Safe Engines
Two major programming languages, Rust (rustc, via its official regex crate) (Foundation, 2024f) and Go (gc) (Team, 2024d), have regex engines that were developed with ReDoS safety in mind. These engines simulate an explicit automaton representation using Thompson’s algorithm (Cox, 2007). They also do not support many E-regex features. They, therefore, offer linear-time match guarantees in input and automaton length.
Although not included in our analysis, independent third-party regex engines like RE2 (Google) (Source, 2024) and Hyperscan (Intel) (Wang et al., 2019; Wang, 2017) are well-known for ReDoS resilience. RE2, a Thompson-style engine, incorporates DFA state caching optimizations (Moseley et al., 2023). While generally ReDoS-safe and non-backtracking, RE2 can, in rare cases, exhibit backtracking behavior (Davis et al., 2021). Similarly, Hyperscan is another efficient, ReDoS-safe engine, utilizing hardware acceleration alongside Glushkov’s NFA construction (Glushkov, 1961). Neither engine supports E-regex features that require backtracking.
Overall, these developments in regex engines demonstrate significant progress in mitigating ReDoS risks but also highlight ongoing challenges in balancing the trade-off between safety and usability. Modern regex engines mitigate ReDoS with non-backtracking algorithms like memoization, Thompson-style simulation, or Brzozowski derivatives to assure linear-time matching. However, the benefit comes at the cost of reduced expressiveness. Non-backtracking engines lack support for many E-regex features. Defenses add timeouts or fallbacks for unsupported E-regexes to maintain codebase compatibility.
@@ 4.2. Updated Measure of ReDoS Vulnerabilities
We answered this question through measurements on a sample from the Davis et al.’s (Davis et al., 2019) polyglot regex corpus, comprising $\sim$500K regexes extracted from open-source repositories across eight programming languages. We used the tool vuln-regex-detector (Davis, 2024b), developed by Davis et al. (Davis et al., 2018a) and based on prior work (Rathnayake and Thielecke, 2017; Wüstholz et al., 2017; Weideman et al., 2016; Shen et al., 2018), to identify potentially vulnerable regexes from the corpus. From its predictions, we randomly sampled 500 regexes labeled as exponential-time candidates (out of $\sim$3K predicted examples) and 500 labeled as polynomial-time candidates (out of $\sim$100K predicted examples). For each sampled regex, we used the same tool to generate corresponding attack input strings.
@@
Our findings are illustrated in Figure 5. For each engine, we depict the number of regexes that exhibited super-linear behavior (as a fraction of the predicted input) in the pre- and post-ReDoS mitigation versions. Not all regexes were supported in all engines, so the candidate pool size is indicated for each engine. Observe that for the engines with mitigations, there is a substantial decrease in the number of viable super-linear regexes—in other words, these mitigations appear to be effective in addressing the common causes of super-linear behavior. Notably, exponential behavior is resolved in C# and JavaScript but persists in Ruby and Java, while polynomial behavior continues to manifest in JavaScript and Java.
Figure 5. Performance of regex matching in old vs. latest engines in different programming languages. The exact engine versions are listed in Appendix D.2.
@@ 5. Discussion
In this section, we distill and contextualize the main findings of our study, integrating insights from our literature review (§3) and engineering review (§4). For each idea, we discuss concrete future work directions for the community to explore.
@@
(1) ReDoS is a peculiar vulnerability: Unlike traditional ones like SQL injections, ReDoS stands out as a contextual issue, more similar to micro-architectural attacks, where exploitability and impact are heavily dependent on the underlying system’s architecture and deployment conditions. For example, a regex pattern vulnerable to catastrophic backtracking in older versions of Node.js is rendered safe in newer versions due to advancements in regex engines (§4). When detecting ReDoS vulnerabilities, it is not enough to analyze the code of a web application, without considering its deployment configurations. For future static analysis work on ReDoS—particularly those targeting multiple languages such as  (Hassan et al., 2023)—we recommend that they consider and model the language-, engine-, and platform-specific details and differences with respect to ReDoS. For example, a super-linear regex might be more serious in a single-threaded web application architecture like Node.js than in inherently parallel ones like Go. For dynamic and hybrid analyses, we recommend a system-level perspective, similar to Shan et al. (Shan et al., 2017), that deploys code under realistic conditions, taking relevant defenses into account.
@@
(2) Strong assumptions about attackers: We find that prior work often makes strong assumptions about attackers (§3.2), such as assuming they can inject arbitrarily large payloads or control the input to any regex. Also, many papers neglect defenses or rely on rules of thumb, like regex matches should not exceed one second (§3.3). We think that future work should challenge these assumptions by performing empirical studies with real systems, including analyses of how many regexes are actually reachable via attacker-controlled inputs. We also need more empirical studies to show the community how modern applications are deployed in the wild—specifically, which configurations they run on and what defenses they employ.
@@
(3) Developers push back: To the best of our knowledge, there has been no public description of a DoS attack that exploited ReDoS, so the risk posed by these vulnerabilities is still unclear, beyond academic measurement studies (Staicu and Pradel, 2018; Barlas et al., 2022). Moreover, only a small fraction of the papers we analyzed (§3) discuss such attacks, often in unrealistic setups. We believe that these unclear risks together with problematic security assumptions are the culprits for ReDoS vulnerabilities being a contentious topic among developers. Snyk has stated that ReDoS is the most neglected vulnerability type (Limited, 2023), implying that developers often ignore such findings. Also, engineers at GitHub consider it a low severity vulnerability, “not particularly serious, but easy to create by accident, obscure to understand, and sometimes tricky to fix” (Backhouse, 2023). More strident voices argue that ReDoS vulnerabilities are “malicious noise” and just lead to security fatigue (Woodruff, 2022). Several ReDoS CVE advisories have been disputed by the maintainers (e.g., CVE-2023-39663, CVE-2022-42969, CVE-2019-11391), indicating disagreement between maintainers and reporters, also reflected by negative comments in GitHub issues (Appendix E).
@@
The blame for this perspective may lie with the cybersecurity academic community. As noted in §3, researchers have performed large-scale measurements of ReDoS as well as produced new and more effective detectors for vulnerable regexes. Their measurement instruments have been integrated into popular security scanners (e.g., the integration of safe-regex into eslint). However, the resulting scanners analyze only the regex, without considering reachability constraints as well as other input constraints. When academics apply these tools, they are likely to filter false positives based on the regex’s context. In contrast, practitioners sometimes make much ado about nothing. This suggests the need for empirical studies of academic tool uptake and more developer-friendly releases. More importantly, future work in the field should perform user studies with developers to better understand their challenges related to ReDoS—such as tool usability, security assumptions, and deployment configurations. In this way, we can better align the community’s assumptions with those of practitioners.
(4) Engines have evolved significantly: It is encouraging that modern language runtimes changed their implementation in response to prior work on ReDoS (§4). These changes were often driven or informed by researchers (e.g., Ruby’s adoption of Davis et al.’s defenses (Davis, 2019; Davis et al., 2021)). We believe that this is a success story of effective collaboration between researchers and practitioners.
@@
Our study shows that modern regex engines have evolved significantly, incorporating advanced techniques such as Thompson’s algorithm, memoization, and resource capping to mitigate ReDoS. However, there are often redundant implementations that are triggered on demand by the engine, when matching a regex. Hence, future work should investigate the deployed fixes to find potential bugs or misbehaving corner cases. We advocate for comparative studies (such as ours in §4.2), where multiple regex engines are evaluated on the same regexes or datasets to systematically assess the strengths and weaknesses of their deployed defenses. The research community should also propose automatic solutions for migrating to updated engines. At the same time, research on ReDoS should take these advancements into account to ensure its findings remain applicable.
@@
(5) Performance bugs vs. security vulnerabilities: Many of our findings may also extend to other algorithmic complexity vulnerabilities. While the threat of weaponizing slow computations is real, in practice, attackers still tend to favor brute-force, volume-based distributed DoS attacks. Researchers should aim to understand why that is the case and, simultaneously, continue to develop defenses against such potential future attacks. We also believe that there are many insufficiently explored, cross-disciplinary opportunities for ideas that aim to bridge the gap between performance, measurement, and security communities. For example, we argue that most performance bugs (e.g., redundant loop traversals (Olivo et al., 2015)) can probably be lifted to a DoS attack when malicious intent is assumed. Thus, future work should further examine the relation between performance bugs and DoS attacks.
@@ 6. Conclusion
In this work, we perform a multi-faceted study of the recent academic work on ReDoS and the new engineering realities in the mainstream programming languages, after initial fixes were deployed in most of them. We report on concrete examples of language runtimes that adopted defenses, which appear to be motivated by academic research. When studying the efficacy of these defenses, we find that they mitigate most ReDoS payloads, but they still leave significant room for attackers to maneuver. By surveying the academic work on ReDoS, we find that many papers in this domain use weak definitions and strong threat models, often considering any slowdown caused by regex matching as a security vulnerability. We advocate for future work to consider the recently-deployed fixes in mainstream programming languages and to evaluate ReDoS attacks under realistic deployment conditions (e.g., against a web application deployed in the cloud using reversed proxies and redundancy). We also propose restricting the attackers’ capabilities to realistic payload sizes and modest bandwidths that comply with the asymmetric DoS setting. More importantly, future work should study the more general problem of weaponizing slow computation, which the attackers might leverage in the near future.
@@ Acknowledgements.
Ethical Considerations
In our judgment, this work presents an acceptable ethical risk-reward tradeoff. Specifically, the literature review (§3) and engineering review (§4) describe and systematize only publicly available resources.
Data Availability
An artifact accompanying this paper is available at https://doi.org/10.5281/zenodo.15515222. It includes: the CVE data analyzed in §2.3.2, the super-linear regex corpus together with the reproduction scripts for the experiments described in §4.2, and the developer-discussion data and GitHub-crawler scripts from Appendix E.
@@ Appendix A Regex Feature Reference
Table 6. Full set of features and notation of regexes, compiled from (Hazel, 1997). $R$ denotes a sub-pattern.
Abbreviation Feature Notation
@@
PNG Named capture group $(?{\textless}\text{name}{\textgreater}R)$
@@
ADD One-or-more repetition $R+$
QST Zero-or-one repetition $R?$
@@
LWB At-least-$m$ repetition $R\{m,\}$
SNG Exactly-$n$ repetition $R\{n\}$
@@
CCC Custom character class $\text{aeiou}]$
RNG Character range $\text{a}-\text{z}]$
NCCC Negated class $\textasciicircum\text{aeiou}]$
ANY Any character .
WSP Whitespace $\setminus\text{s}$
DEC Numeric $\setminus\text{d}$
WRD Word $\setminus\text{w}$
NWSP Non-whitespace $\setminus\text{S}$
NDEC Non-numeric $\setminus\text{D}$
NWRD Non-word $\setminus\text{W}$
VWSP Vertical space $\setminus\text{v}$
@@
STR Start-of-string/line $\textasciicircum R,\;\setminus\text{A}R$
@@
WNW Word/non-word boundary $\setminus\text{b}$
NWNW Negated WNW boundary $\setminus\text{B}$
PLA Positive lookahead $(?=R)$
NLA Negative look-ahead $(?!R)$
PLB Positive lookbehind $(?{\textless}=R)$
NLB Negative look-behind $(?{\textless}!R)$
@@
BKR Numeric backreference $(R)\ldots\setminus 1$
BKRN Named backreference $(?{\textless}\text{name}{\textgreater}R)\ldots\setminus\text{k{\textless}name{\textgreater}}$
@@
LZY Non-greedy repetition $R*?,\;R+?,\;R\{m,n\}?$
ATM Atomic group $(?{\textgreater}R)$
POS Possessive quantifier R++,R+R++,\;R*+
Table 7. Examples of excluded works with explanations.
Work Reason for Exclusion
Cox (Cox, 2007) Evaluation does not address ReDoS-specific scenarios.
Mamouras et al. (Mamouras and Chattopadhyay, 2024) Evaluation does not address ReDoS-specific scenarios.
Shan et al. (Shan et al., 2017) Addresses general DoS attacks.
Wei et al. (Meng et al., 2018) Addresses general DoS attacks.
Table 8. Regex engine implementations selected for evaluating ReDoS defenses.
JavaScript Node.js—V8 (Foundation, 2024c; Project, 2024b) Deno, Bun, SpiderMonkey, JavaScriptCore Dominates server-side JavaScript usage (98.8% market share) (Q-Success, 2024)
Ruby MRI/CRuby (Community, 2024c) JRuby, Rubinius, mruby, RubyMotion Reference implementation (Community, 2024b)
C# .NET (Foundation, 2024a) Mono Reference implementation (Contributors, 2024)
Java OpenJDK (Corporation, 2024) Amazon Corretto, Azul Zulu, Liberica JDK, Eclipse Adoptium… Most widely used JDK of 2024 with a 20.8% usage rate (New Relic, 2024)
PHP Zend Engine (Group, 2024a) HHVM, PeachPie, Quercus, Parrot Standard PHP interpreter is driven by the Zend Engine (Group, 2024b, c)
Perl perl5 (Perl.org, 2024b) N/A (for Perl 5) Only implementation for Perl 5 (Perl.org, 2024a)
Rust rustc (Foundation, 2024g) mrustc, gccrs Official and only fully functional compiler for Rust (Foundation, 2024h)
Go gc (Team, 2024b) gccgo, gofrontend, TinyGo, GopherJS, yaegi Official compiler included in Go releases (Team, 2024c)
Python CPython (Foundation, 2024e) PyPy, Stackless Python, MicroPython, CircuitPython, IronPython, Jython Reference implementation (Foundation, 2024d)
PL: Programming Language, Impl.: Implementation (runtime environment, engine, interpreter, or compiler)
Table 9. Versions used in the ReDoS defense measurements.
Language Old Version New Version
JavaScript (Node.js—V8) v15.14.0 v22.2.0
Ruby (MRI/CRuby) 3.1.6 3.3.2
C# (.NET) 6.0.420 7.0.407
Java (OpenJDK) 8u342 23
PHP (Zend Engine) 5.6.40 8.3.7
Perl (perl5) 5.8.14 5.38.2
Rust (rustc) 1.12.1 1.78.0
Go (gc) 1.5.4 1.22.4
Python (CPython) 3.6.15 3.12.3